. A spherical shaped vessel of 1.4 m diameter is 90 mm thick. Find the rate

of heat leakage, if the temperature difference between the inner and outer surfaces is 220°C.

Thermal conductivity of the material of the sphere is 0.083 W/m°C.


1
Expert's answer
2021-09-03T08:25:02-0400

Solution.

"r_2=0.7m; r_1=0.7-0.09=0.61m;"

"dT=220^oC; k=0.083W\/mK;"

"Q=\dfrac{dT}{R};"

"R=\dfrac{r_2-r_1}{4\pi kr_1r_2}=\dfrac{0.7-0.61}{4\sdot3.14\sdot0.083\sdot0.61\sdot0.7}=0.2022m;"

"Q=\dfrac{220}{0.2022}=1088W/mK=1.088kW/mK;"

Answer: "Q=1.088kW\/mK."